OpenStudy (abbles):

Given matrices A, B, and C below, perform the indicated operations if possible. Precalc question?

OpenStudy (abbles):

If the operation is not possible, explain why.

OpenStudy (abbles):

A = 2 -1 0 0 5 0.3 1 4 10 B = 5 0 2 1 -3 9 2 0 4 C = [1 3 5]

jimthompson5910 (jim_thompson5910):

You posted `perform the indicated operations if possible` but you forgot to post the actual operations.

OpenStudy (abbles):

Yeah, I figured it out elsewhere. Thanks anyway!

OpenStudy (abbles):

Just reviewing the unit (all the questions I missed in the past) for an upcoming test.